OpenStudy (hba):

In the draw of a card from a deck of ordinary playing cards,find the probability that the card is a 7 or a club ?

OpenStudy (hba):

@hartnn

hartnn (hartnn):

how many 7 are there in a deck of ordinary playing cards?

OpenStudy (hba):

4

OpenStudy (hba):

I GUESS

hartnn (hartnn):

yup.

hartnn (hartnn):

how many clubs are there ?

OpenStudy (hba):

13

OpenStudy (hba):

I GUESS

hartnn (hartnn):

right.

hartnn (hartnn):

and total cards ?

OpenStudy (zarkon):

you 'GUESS' a lot :)

OpenStudy (hba):

13*4=52+2 JOKERS =54 I GUESS @Zarkon :D

OpenStudy (zarkon):

usually don't include jokers

hartnn (hartnn):

don't count joker.

OpenStudy (hba):

Ok 52

hartnn (hartnn):

so probability is defined as favourable / total so, can u tell me the probability of getting 7 ?

OpenStudy (hba):

4/52

hartnn (hartnn):

right, same way find, probability of setting a club. so now you let P(S) = probability of getting 7 let P(C) = probability of getting club you want P(S OR P) u know the formula ?

hartnn (hartnn):

*getting

OpenStudy (hba):

Actually we will have to count it as 12 now because we dont want to repeat the 7 because its already been taken if i am not wrong and i dont really know how to do this so it would be better to guide me to the answer :)

hartnn (hartnn):

ok, how many 7 'and' club cards are there ?

OpenStudy (hba):

In total 16

OpenStudy (hba):

because we are not repeating the 7 of clubs

hartnn (hartnn):

P(S OR C) = P(S)+P(C)-P(S and C) thats why i asked you 7 'and' club cards.

OpenStudy (hba):

Ok Using this 13+4-1=16

hartnn (hartnn):

ok, those are probabilities, so u get 13/52 +4/52- 1/52 ok? which is 16/52 = ?

OpenStudy (hba):

ok

hartnn (hartnn):

so u get P(card is a 7 or a club) = 16/52 = ?

OpenStudy (hba):

4/13

OpenStudy (hba):

is the answer

hartnn (hartnn):

yup

OpenStudy (hba):

Thanks :)